OpenStudy (anonymous):

Find an exact value. cos(-7pi/12)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

cos(-7pi/12)=cos(7pi/12)=cos(105)

OpenStudy (anonymous):

cos(105)=-cos(75)

OpenStudy (mathmale):

First, a question for you: What two simple fractions with denominators 3 and 4 respectively add up to 7/12?

OpenStudy (anonymous):

so it would be C? @mustafa2014

OpenStudy (anonymous):

3/3 4/4 @mathmale

OpenStudy (anonymous):

@mustafa2014

OpenStudy (mathmale):

3/3 and 4/4 add up to 2, so that could not be the correct combination of fractions.

OpenStudy (mathmale):

My question is very relevant to solving this trig problem, in case you're wondering.

OpenStudy (anonymous):

@mustafa2014 ?????????

OpenStudy (anonymous):